40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 253 of 341  Not logged in ELOG logo
ID Date Author Type Category Subject
  4448   Mon Mar 28 16:24:35 2011 kiwamuUpdateGreen Lockingpower budget on PSL table

   I measured some laser powers associated with the beat-note detection system on the PSL table.

The diagram below is a summary of the measurement. All the data were taken by the Newport power meter.

 The reflection from the beat-note PD is indeed significant as we have seen.

In addition to it the BS has a funny R/T ratio maybe because we are using an unknown BS from the Drever cabinet. I will replace it by a right BS.

RFPD.png

(background)

 During my work for making a noise budget I noticed that we haven't carefully characterize the beat-note detection system.

The final goal of this work is to draw noise curves for all the possible noise sources in one plot.

To draw the shot noise as well as the PD dark noise in the plot, I started collecting the data associated with the beat-note detection system.

 

(Next actions)

 * Estimation and measurement of the shot noise

 * measurement of the PD electrical noise (dark noise)

 * modeling for the PD electrical noise

 * measurement of the doubling efficiency

 * measurement of an amplitude noise coupling in the frequency discriminators

  4447   Mon Mar 28 16:19:23 2011 steveFrogsPhotosvisithing 5th graders

Suresh is captivating his audience with gravity waves on last Friday, March 25

Attachment 1: P1070475.JPG
P1070475.JPG
  4446   Mon Mar 28 15:49:18 2011 josephbUpdateCDSLessons from LST

[Koji,Joe]

PART 1:

Koji was unable to build his c1lst model first thing this morning.  Turns out there was  a bug with RCG parser that was introduced on Friday when we did the RCG updates.  We talked Alex who did a quick comment fix.  The diff is as follows:

Index: Parser3.pm
==============================

=====================================
--- Parser3.pm  (revision 2328)
+++ Parser3.pm  (working copy)
@@ -1124,8 +1124,8 @@
  print "Flattening the model\n";
  flatten_nested_subsystems($
root);
  print "Finished flattening the model\n";
-  CDS::Tree::do_on_nodes($root, \&remove_tags, 0, $root);
-  print "Removed Tags\n";
+  #CDS::Tree::do_on_nodes($root, \&remove_tags, 0, $root);
+  #print "Removed Tags\n";
  #print "TREE\n";
  #CDS::Tree::print_tree($root);
  CDS::Tree::do_on_nodes($root, \&remove_busses, 0, $root);

This was some code to remove TAGs from the .mdl file for some reason which I do not understand at this time.  I will ask tommorrow in person so I can understand the full story.

PART 2:

Koji then rebuilt and started the c1lst process.  This is his new test version of the LSC code.  We descovered (again) that when you activate too many DAQ channels (simply uncommenting them, not even recording them with activate=1 in the .ini file) that the frame builder crashes.  In addition, the c1lsc machine, which the code was running on, also hard crashed.

When a channel gets added to the .ini file (or uncommented) it is sent to the framebuilder, irregardless of whether its recorded or not by the frame builder.  There is only about 2 megabytes per second bandwidth per computer.  In this case we were trying to do something like 200 channels * 16384 Hz * 4 bytes = 13 megabytes per second.

The maximium number of 16384 channels is roughly 30, with little to no room for anything else.  In addition, test points use the same allocated memory structure, so that if you use up all the capacity with channels, you won't be able to use testpoints to that computer (or thats what Alex has led me to believe).

The daqd process then core dumped and was causing all sorts of martian network slowdowns.  At the same time, the c1lsc computer crashed hard, and all of the front end processes except for the IOP on c1sus crashed.

We rebooted c1lsc, and restarted the c1sus processes using the startc1SYS scripts.  However, the c1susfe.ko apparently got stuck in a wierd state.  We were completely unable to damp the optics and were in general ringing them up severely.  We tried debugging, including several burt restores and single path checks.

Eventually we decided to reboot the c1sus machine after a bit of debugging.  After doing a burt restore after the reboot, everything started to damp and work happily.  My best guess is the kernel module crashed in a bad way and remained in memory when we simply did the restart scripts.

 

  4445   Mon Mar 28 15:18:04 2011 josephbUpdateCDSCDS updates on Friday

Last Friday, we discovered a bug in the RCG where the delay part was not actually delaying.  We reported this to Alex who promptly put a fix in the same day.  This allowed Matt's newly proposed frequency discriminator to work properly.

It also required a checkout of the latest RCG code (revision 2328), and rebuild of the various codes.  We backed up all the kernel and executables first such as mbuf.ko and awgtpman.

We did the following:

1) Log into the fb machine.

2) Go to /opt/rtcds/caltech/c1/core/advLigoRTS/src/drv/mbuf and run make.  Copy the newly built mbuf.ko file to /diskless/root/modules/2.6.34.1/kernel/drivers/mbuf/mbuf.ko on the fb machine.

3) Use "sudo cp" to copy the newly built mbuf.ko file to /diskless/root/modules/2.6.34.1/kernel/drivers/mbuf/

4) Go to /cvs/cds/rtcds/caltech/c1/core/advLigoRTS/src/gds and run make.

5) Copy the newly built awgtpman executable to /opt/rtcds/caltech/c1/target/gds/bin/

6) Go to /opt/rtcds/caltech/c1/core/advLigoRTS/src/mx_stream/ and run make.

7) Copy the newly built mx_stream executable to /opt/rtcds/caltech/c1/target/fb/

  4444   Fri Mar 25 11:16:19 2011 josephbFrogsGreen Lockingdigital frequency counting

I modified the c1gfd.mdl simulink model.  I made a backup as c1gfd_20110325.mdl. 

The first change was to use a top_names block to put everything in.  The block is labeled ALS.  So all the channels will now be C1:ALS-GFD_SOMETHING.  This means medm channel names will need to be updated.  Also, the filter modules need to be updated in foton because of this.

I then proceeded to add the suggested changes made by Matt.  To avoid a divide by zero case, I added a saturation part which saturates at 1e-9 (note this is positive) and 1e9.

Quote:

Today we tried the Schmitt trigger DFD, and while it works it does not improve the noise performance.  At least part of our problem is coming from the discrete nature of our DFD algorithm, so I would propose that an industrious day job person codes up a new DFD which avoids switching.  We can probably do this by mixing the input signal (after high-passing) with a time-delayed copy of itself... as we do now, but without the comparator.  This has the disadvantage of giving an amplitude dependent output, but since we are working in the digital land we can DIVIDE.  If we mix the signal with itself (without delay) to get a rectified version, and low-pass it a little, we can use this for normalization.  The net result should be something like:

output = LP2[ s(t) * s(t - dt) / LP1[ s(t) * s(t) ]],

where s(t) is the high-passed input and LP is a low-pass filter.  Remember not to divide by zero.

 

 

Attachment 1: C1GFD.png
C1GFD.png
  4443   Fri Mar 25 08:58:38 2011 AidanUpdateTreasureCleared stuff off SP table

I tidied up some of the stuff that was on the SP table. The ISS box that has been sitting on there for months is now underneath the X-arm on top of the spare Marconi which is stored there.

 

 

Attachment 1: ISS_box.JPG
ISS_box.JPG
  4442   Fri Mar 25 01:27:29 2011 mevansFrogsGreen Lockingdigital frequency counting

Today we tried the Schmitt trigger DFD, and while it works it does not improve the noise performance.  At least part of our problem is coming from the discrete nature of our DFD algorithm, so I would propose that an industrious day job person codes up a new DFD which avoids switching.  We can probably do this by mixing the input signal (after high-passing) with a time-delayed copy of itself... as we do now, but without the comparator.  This has the disadvantage of giving an amplitude dependent output, but since we are working in the digital land we can DIVIDE.  If we mix the signal with itself (without delay) to get a rectified version, and low-pass it a little, we can use this for normalization.  The net result should be something like:

output = LP2[ s(t) * s(t - dt) / LP1[ s(t) * s(t) ]],

where s(t) is the high-passed input and LP is a low-pass filter.  Remember not to divide by zero.

 

  4441   Thu Mar 24 19:48:13 2011 Aidan, KiwamuUpdateGreen LockingDesigns for permanent electronics for ALS

Kiwamu and I looked at all the electronics that are currently in place for the green locking on the X-arm and have made a set of block diagrams of the rack mounted units that we should build to replace the existing ... "works of art" that sprawl around out there at the moment.

Main items

1. "ETM Green Oscillator/PDH support box". Not a great name but this would provide the local oscillator signal for the end PDH (with a controllable phase rotator) as well as the drive oscillator for the end laser PZT. Since we need to hit a frequency of 216.075kHz with a precision that Kiwamu needs to determine, we'd need to be able to tune the oscillator ... it needs to be a VCO. It'd be nice to be able to measure the output frequency so I've suggested dividing it down by N times to put it into the DAQ - maybe N = 2^7 = 128x to give a measured frequency of around 1.7kHz. Additionally this unit will sum the PDH control signal into the oscillation. This box would support the Universal PDH box that is currently at the X-end.

2. "Vertex X-arm beatnote box" - this basically takes the RF and DC signals from the beatnote PD and amplifies them. It provides a monitor for the RF signal and then converts the RF signal into a square wave in the comparator.

3. "Mixer Frequency Discriminator" - just the standard MFD setup stored in a box. For temperature stability reasons, we want to be careful about where we store this box and what it is made of. That's also the reason that this stage is separated from the X-arm beatnote box with it's high-power amps.

Other things

4. RS232 and EPICS control of the doubling ovens

5. Intensity stabilization of the End Laser

P.S. I used Google Diagrams for the pictures.

Attachment 1: GreenLockingElectronics.pdf
GreenLockingElectronics.pdf
Attachment 2: GreenEndPDHsupportboxandLO.pdf
GreenEndPDHsupportboxandLO.pdf
Attachment 3: VertexBeatnoteAmplifierandComparator.pdf
VertexBeatnoteAmplifierandComparator.pdf
Attachment 4: MixerFrequencyDiscriminator.pdf
MixerFrequencyDiscriminator.pdf
  4440   Thu Mar 24 16:33:32 2011 BryanConfigurationGreen LockingPSL vs Y arm laser temperature pairing

X_arm and Y_arm vs PSL comparison.

 

 

Just a quick check of the performance of the X arm and Y arm lasers in comparison to the PSL. Plotting the data from the X arm vs PSL and Y arm vs PSL on the same plot shows that the X arm vs the PSL has no observable trending of mode-hopping in the laser, while the Y arm vs the PSL does. Suspect this is due to the fact that the X arm and PSL are both Innolight lasers with essentially identical geometry and crystals and they'll tend to mode-hop at roughly the same temperatures - note that the Xarm data is rough grained resolution so it's likely that any mode-hop transitions have been skipped over. The Lightwave on the other hand is a very different beast and has a different response, so won't hop modes at the same temperatures.

Given how close the PSL is to one of the mode-transition regions where it's currently operating (32.26 degC) it might be worth considering shifting the operating temperature down one degree or so to around 31 degC? Just to give a bit more headroom. Certainly worth bearing in mind if problems are noticed in the future.

PSL_vs_X-end_T_data_fine.jpg

  4439   Thu Mar 24 15:30:59 2011 BryanConfigurationGreen LockingPSL vs Y arm laser temperature pairing
Fine-grained temperature vs temperature data around the current operating point of the PSL laser.
 
The last set of data was taken in 1 degreeC steps, but we want a bit more detail to find out what happens around the current PSL operating point. So we took some data with a 0.1 degC resolution.

The good news is that we seem to be running in a linear region of the PSL laser with a degree or so of range before the PSL Innolight laser starts to run multi-mode. On the attached graph we are currently running the PSL at 32.26degrees (measured) which puts us in the lower left corner of the plot. The blue data is the Lightwave set temperature (taken from the display on the laser controller) and the red data is the Lightwave laser crystal measured temperature (taken from the 10V/degC calibrated diagnostic output on the back of the laser controller - between pins 2 and 4).

The other good news is that we can see the transition between the PSL laser running in one mode and running in the next mode along. The transition region has no data points because the PMC has trouble locking on the multi-mode laser output - you can tell when this is happening because, as we approach the transition the PMC transmitted power starts to drop off and comes back up again once we're into the next mode region (top left portion of the plot).

 

The fitted lines for the region we're operating in are:

Y_arm_Temp_meas = 0.95152*T_PSL + 3.8672

Y_arm_Temp_set = 0.87326*T_PSL + 6.9825

Temp_Beating_Day02_html_m2d719182.jpg

 

  4438   Thu Mar 24 13:56:05 2011 josephbUpdateelogelog restarted at 1:55pm

Restarted elog.

  4437   Thu Mar 24 13:50:30 2011 BryanConfigurationGreen LockingY arm laser

 Just a quick update... the Lightwave laser has now been moved up to the end of the Y arm. It's also been mounted on the new mounting block and heatsinks attached with indium as the heat transfer medium.

A couple of nice piccies...IMG_0188.JPG

Attachment 2: IMG_0190.JPG
IMG_0190.JPG
  4436   Thu Mar 24 01:16:19 2011 SureshSummaryGreen LockingY-END green equipment is all available
There was a 2" mirror mount among the spares on the PSL table.  It has a window LW-3-2050 UV mounted in it.  I
have moved it to the Y-end table.  We seem to have run out of 2" mirror mounts ...
  4435   Wed Mar 23 19:16:17 2011 AidanSummaryGreen LockingY-END green equipment is all available

With the exception of a 2" mirror mount, I've confirmed that we have everything for the Y-end green production and mode-matching.

We need to calculate a mode-matching solution for the Lightwave laser so that it gives the correct beam size in the doubling crystal.

Additionally, Rana has suggested that we change the pedestals from the normal 1" diameter pedestal+fork combo  to the 3/4" diameter posts and wider bases that are used on the PSL table (as shown in the attached image).

Attachment 1: three-quarter_inch_pedestal.jpg
three-quarter_inch_pedestal.jpg
  4434   Wed Mar 23 16:06:20 2011 Larisa ThorneUpdateElectronicsUpdate on cable laying

 [Steve, Suresh, Larisa]

The following cables were laid today: ETMYT, ETMY, IFOPO, MC1, OMCR, AS Spare, and MC2T.

 

Though the paper suggested 135' for the MC2T, we used a 110'. This is too short: need at least another 15' for the MC2T.

The RCR cable wasn't crossed off on the list, but a cable exists at the RCR cable which is black and is labeled (old label, 75 ohms)

There was no indication of which length was needed for MC1, so a 95' cable was used.

  4433   Wed Mar 23 14:19:35 2011 KojiSummaryGeneralGrand Plan

This is the grand plan we talked about in the beginning of the meeting.

  • (Kiwamu) X-end Green cleaning up / Prep for DRMI
  • (Bryan) Y-end Green
  • (Suresh) Help Bryan / RF (w. Kevin)
  • (Jenne) MC WFS / Y-arm IR alignment / MC adaptive feedforward (incl. CDS)
  • (Koji) LSC
  • (Joe) CDS cleaning up
  • (Jamie) Help Joe / Noise Budget
  • (Larisa) PMC scan / PSL photo&diagram
  • (Barbarela) ASS
  4432   Wed Mar 23 12:40:22 2011 Chief Recycling OfficerHowToEnvironmentRecycle stuff!

The following is a message from the LIGO 40m Chief Recycling Officer:

Please get up off your (Alignment Stabilization Servo)es and recycle your bottles and cans!  There is a recycling bin in the control room.  Recent weeks have seen an increase in number of bottles/cans thrown away in the regular garbage.  This is not cool. 

Thank you,

---L4mCRO

  4431   Wed Mar 23 10:34:17 2011 josephbUpdateCDSTrend issue fixed

[Joe, Alex]

Yesterday during the day, Alex ran a script to fix the time stamps in the trends files we had messed up back during the daqd change overs around Feb 17th and 23rd.  See this elog for more information on the trend problem.

Due to how the script runs, basically taking all the data and making a new copy with the correct time stamps, the data collected while the script was running didn't get converted over.  So when he did the final copy of the corrected data, it created a several hour gap in the data from yesterday during the day time.

The original files still exist on the fb machine in /frames/trend/minute_raw_22mar2011 directory.

 

  4430   Wed Mar 23 09:54:46 2011 steveOmnistructurePhotosLSC visitors

The 40m lab was visited by  ~ 30 LSC members  the end of last week.

Attachment 1: P1070467.JPG
P1070467.JPG
Attachment 2: P1000414.jpg
P1000414.jpg
  4429   Wed Mar 23 09:48:20 2011 steveUpdatePSLPSL enclosure: gets new window & laser

Solid door, numbered 4 at south west corner of PSL enclosure was replaced by laser protective window.

The carpenter shop's Mark is making 4 more identical ones for the east side.

 

The Lightwave NPRO126 of 700mW was moved from the AP-table into the PSL-enclosure temporarily.

It's emergency shutdown switch can be seen at the center bottom picture

Attachment 1: P1070471.JPG
P1070471.JPG
  4428   Wed Mar 23 08:50:36 2011 AidanUpdateGreen Lockingservo handig off

Nicely done!

 

Quote:

Succeeded in handing off the servo from the green to the red.

stripTransit_edit.png

 

 

 

Attachment 1: green-to-red.jpg
green-to-red.jpg
  4427   Wed Mar 23 05:11:08 2011 kiwamuUpdateGreen Lockingservo handig off

Succeeded in handing off the servo from the green to the red.

stripTransit_edit.png

 

 


(noise performance)

 This time we found that the fluctuation in the IR signals became lesser as the gain of the ALS servo increased.

Therefore I increased the UGF from 40 Hz to 180 Hz to have less noise in the IR PDH signal.

Here is a preliminary plot for today's noise spectra.

noise_curve20110322_hUGF.png

The blue curve is the ALS in-loop spectrum, that corresponds to the beat fluctuation.

The red curve is an out-of-loop spectrum taken by measuring the IR PDH signal.

Since the UGF is at about 180 Hz the rms is integrated from 200 Hz.

The residual displacement noise in the IR PDH signal is now 1.2 kHz in rms.

I am going to analyze this residual noise by comparing with the differential noise that I took yesterday (see the last entry ).

  4426   Wed Mar 23 00:51:47 2011 kiwamuUpdateGreen Lockingplan for tomorrow

  - Plan for tomorrow

    * Video cable session (I need ETMY_TRNAS) (team)

    * Characterization of the Y end laser  (Bryan / Suresh)

    * LPF for the X end laser temperature control (Larisa)

    * Frequency Divider  (Matt)

    * X end mechanical shutter (Kiwamu)

  4425   Tue Mar 22 19:03:45 2011 BryanConfigurationGreen LockingPSL vs Y arm laser temperature pairing

Quote:

 I'm going to take the easy question - What are the pink data points??

And I'm going to answer the easy question - they're additional beat frequency temperature pair positions which seem to correspond to additional lines of beat frequencies other than the three highlighted, but that we didn't feel we had enough data points to make it worthwhile fitting a curve.

It's still not entirely clear where the multiple lines come from though - we think they're due to the lasers starting to run multi-mode, but still need a bit of thought on that one to be sure...

  4424   Tue Mar 22 16:39:51 2011 kiwamuUpdateGreen Lockingcomaprator installed : 80 pm residual displacement

 A comparator has been installed before the MFDs (mixer-based frequency discriminator) to eliminate the effect from the amplitude fluctuation (i.e. intensity noise).

As a result we reached an rms displacement of 580 Hz or 80 pm.

 

(differential noise measurement)

noise_curve20110322.png

  Here is the resultant plot of the usual differential noise measurement.

The measurement has been done when the both green and red lasers were locked to the X arm.

In the blue curve I used only MFD. In the black curve I used the combination of the comparator and the MFD.

Noise below 3 Hz become lower by a factor of about 4, resulting in a better rms integrated from 40 Hz.

Note that the blue and the black curve were taken while I kept the same lock.

A calibration was done by injecting a peak at 311 Hz with an amplitude of 200 cnt on the ETMX_SUS_POS path.

 

(installation)

  Yesterday Koji modified his comparator circuit such that we can take a signal after it goes thorough the comparator.

The function of this comparator is to convert a sinusoidal signal to a square wave signal so that the amplitude fluctuation doesn't affect the frequency detection in the MFD.

I installed it and put the beat-note signal to it. Then the output signal from the comparator box is connected to the MFDs.

The input power for the comparator circuit has been reduced to -5 dBm so that it doesn't exceeds the maximum power rate.

  4423   Tue Mar 22 00:23:20 2011 JenneConfigurationGreen LockingPSL vs Y arm laser temperature pairing

Quote:

 OK. Today we did the same type of measurement for the Y arm laser as was done for the X arm laser here: http://nodus.ligo.caltech.edu:8080/40m/3759 

And attached here is a preliminary plot of the outcome - oddities with adding on the fitted equations, but they go as follows

(Red)    T_yarm = 1.4435*T_PSL - 14.6222

(Blue)    T_yarm = 1.4223*T_PSL - 10.9818

(Green) T_yarm = 1.3719*T_PSL - 6.3917

 

 It's a bit of a messy plot - should tidy it up later...

 I'm going to take the easy question - What are the pink data points??

  4422   Tue Mar 22 00:03:29 2011 BryanConfigurationGreen LockingPSL vs Y arm laser temperature pairing

 OK. Today we did the same type of measurement for the Y arm laser as was done for the X arm laser here: http://nodus.ligo.caltech.edu:8080/40m/3759 

And attached here is a preliminary plot of the outcome - oddities with adding on the fitted equations, but they go as follows

(Red)    T_yarm = 1.4435*T_PSL - 14.6222

(Blue)    T_yarm = 1.4223*T_PSL - 10.9818

(Green) T_yarm = 1.3719*T_PSL - 6.3917


 

 PSL_vs_Y_arm_Temperatures.png

It's a bit of a messy plot - should tidy it up later...

  4421   Tue Mar 22 00:01:25 2011 kiwamuSummaryGreen Lockingplan for daytime tasks

Some tasks for the daytime tomorrow.

  * Beam profile measurements of the Y end laser  (Suresh / Bryan)

  * Taking care of CDS and the simulated plant (Jamie / Joe)

  * Reconnect the X end mechanical shutter to 1X9 (Kiwamu)

  * LPF for the X end temperature feedback (Larisa)

  4420   Mon Mar 21 18:34:10 2011 kiwamuUpdateGreen Lockingadded a new ADC channel on 1X9

I added a new ADC channel for a DC signal from the X end green PD.

It is called C1:GCX-REFL_DC and connected to adc_0_1, which is the second channel of ADC_0.

 

By the way, when I tried connecting it to an ADC I found that most of the channels on the AA board on 1X9 were not working.

Since the outputs form the board are too small the circuits may have benn broken. See the picture below.

In addition to that  I realized that the signal from the PDH box for the temperature actuation is limited by +/- 2V due to the range of this AA board.

In fact the signal is frequently saturated due to this small voltage range.

We have to enlarge the range of this AA board like Valera did before for the suspensions (see this entry).

aa_board_1X9.jpg

  4419   Mon Mar 21 16:49:11 2011 kiwamuSummaryGreen Lockingplan for this week

- Plan for this week

  * Intensity stabilization for the end green laser (Matt / Kiwamu)

  * Hand off the servo from Green to Red (Matt / Kiwamu)

  * Y end green locking (Suresh / Bryan) (rough schedule)

  * Reconnect the X end mechanical shutter to 1X9 (Kiwamu)

  * Connect the end DCPD signal to a DAC (done)

  * Make a LPF in a Pomona box for the temperature (Larisa)

  * Clean up and finalize the X end setup (Kiwamu)

  * Make a item lists for electronics. Order the electronics. (Aidan / Kiwamu)

  4418   Mon Mar 21 16:31:02 2011 steveUpdateSAFETYBryan Barr received safety training

Bryan Barr is visiting us from Glasgow for a month. He received 40m specific safety training on Friday.

  4417   Mon Mar 21 13:26:25 2011 KojiUpdatePSLPMC Trans/RFPDDC

PMC TRANS/REFL on MEDM showed red values for long time.
TRANS (a.k.a C1:PSL-PSL_TRANSPD) was the issue of the EPICS db.

REFL (a.k.a. C1:PSL-PMC_RFPDDC) was not physically connected.
There was an unknown BNC connected to the PMC DC output instead of dedicated SMA cable.
So they were swapped.

Now I run the following commands to change the EPICS thresholds:

ezcawrite C1:PSL-PMC_PMCTRANSPD.LOLO 0.8
ezcawrite C1:PSL-PMC_PMCTRANSPD.LOW 0.85
ezcawrite C1:PSL-PMC_PMCTRANSPD.HIGH 0.95
ezcawrite C1:PSL-PMC_PMCTRANSPD.HIHI 1

ezcawrite C1:PSL-PMC_RFPDDC.HIHI 0.05
ezcawrite C1:PSL-PMC_RFPDDC.HIGH 0.03
ezcawrite C1:PSL-PMC_RFPDDC.LOW 0.0
ezcawrite C1:PSL-PMC_RFPDDC.LOLO 0.0

As these commands only give us the tempolary fix, /cvs/cds/caltech/target/c1psl/psl.db was accordingly modified for the permanent one.

grecord(ai,"C1:PSL-PMC_RFPDDC")
{
        field(DESC,"RFPDDC- RFPD DC output")
        field(DISV,"1")
        field(SCAN,".1 second")
        field(DTYP,"VMIVME-3113")
        field(INP,"#C0 S32 @")
        field(EGUF,"10")
        field(EGUL,"-10")
        field(EGU,"Volts")
        field(PREC,"3")
        field(LOPR,"-10")
        field(HOPR,"10")
        field(AOFF,"0")
        field(LINR,"LINEAR")
        field(LOW,"0.0")
        field(LSV,"MINOR")
        field(LOLO,"0.0")
        field(LLSV,"MAJOR")
        field(HIGH,"0.03")
        field(HSV,"MINOR")
        field(HIHI,"0.05")
        field(HSV,"MAJOR")
}

grecord(ai,"C1:PSL-PMC_PMCTRANSPD")
{
        field(DESC,"PMCTRANSPD- pre-modecleaner transmitted light")
        field(DISV,"1")
        field(SCAN,".1 second")
        field(DTYP,"VMIVME-3123")
        field(INP,"#C0 S10 @")
        field(EGUF,"10")
        field(EGUL,"-10")
        field(EGU,"volts")
        field(PREC,"3")
        field(LINR,"LINEAR")
        field(HOPR,"10")
        field(LOPR,"-10")
        field(AOFF,"0")
        field(LOW,"0.8")
        field(LSV,"MINOR")
        field(LOLO,"0.85")
        field(LLSV,"MAJOR")
        field(HIGH,"0.95")
        field(HSV,"MINOR")
        field(HIHI,"1.00")
        field(HSV,"MAJOR")
}

  4416   Fri Mar 18 17:55:58 2011 SureshConfigurationGreen LockingWork Plan for Y-end Aux laser installation

A rough time-table and the various tasks are given below:

Note:  700mW NPRO sitting on AP table (Model No: 126-1064-700, Sl No. 415)  = Alberto's laser

 

 

Y-arm Aux laser installation
 1

Temperature dependence of frequency of Alberto's laser:

 a) Shifting Alberto's Laser (AL) to the PSL table and setting up a beat frequency measurement between AL and PSL

 b) Determining the frequency vs Temperature curve for the AL

Mar 21st to 25th Bryan and Suresh
2 Re-positioning the Input beam onto the IP-ANG-PD and realigning the X-arm Mar 21st to 25th Kiwamu and his 'team'  :-)
3

Repositioning the optics on the Y-end  table and relocating Alberto's laser ( at this point it will be rechiristened as Y-End-NPRO )

Mar 27th - 28th
Bryan and Suresh
4 Maximising the doubling effiiciency and obtaining the PD and QPD signals into the CDS Mar 29th - Apr 1st "
5 Aligning the Y-end green to pass through the Y-arm and locking the green to the Y arm Apr 3 - 8th "
6 Aligning the IR beam to the Y- arm and locking the Y arm to the IR Apr 10 - 15 "

 

  4415   Fri Mar 18 17:25:21 2011 josephbUpdateCDSLockins in c1sus update, suspension screens updated

I updated our lockin simulink pieces to use the newer, more streamlined lockin piece that is currently in CDS_PARTS (with new documentation block!).  It means we are no longer passing clock signals through three levels of boxes.

In order to use the piece, you need to right click on it after copying from CDS_PARTS and go to Link Options->Disable Link.  This forces the .mdl to save all the relevant information about the block rather than just a pointer to the library.  I talked with Rolf and Alex today and we discussed setting up another model file, non-library format for putting generically useful user blocks into, rather than using the CDS_PARTS library .mdl.

The BS, ITMX, ITMY, PRM, SRM, ETMX, ETMY now have working lockins, with the input matrix to them having the 2nd input coming from LSC_IN, the 3rd from the oplev pitch, and the 4th from oplev yaw.

This necessitated a few name changes in the medm screens.  I also changed the lockin clock on/off switch to a direct amplitude entry, which turns green when a non-zero value is entered.

Currently, the Mode cleaner optic suspension screens have white lockins on them.  I started modifying a new set of screens just for them, and will modify the generate_master_screens.  Unfortunately, this requires modifying two sets of suspension screens going forward - the main interferometer optics and the MC optics.

  4414   Fri Mar 18 16:31:11 2011 Suresh UpdateGreen LockingRe: Y arm plan for today

The reason for using Alberto's laser is that some amount of work has already gone into characterising its phase noise.  Ref elog entry 2788

  4413   Fri Mar 18 16:06:30 2011 kiwamuUpdateGreen LockingRe: Y arm plan for today

We use Alberto's laser for the Y end Green Locking.

Quote:

 Which laser are we going to use,  Alberto's laser or MOPA laser ?

 

  4412   Fri Mar 18 14:18:00 2011 kiwamuSummaryGeneralnew laser pointers

Just for a record. We got 4 new laser pointers (2 greens, 1 blue, and 1 green and red combination). Don't lose them.

They reside in a bucket on the SP table, where IR viewers and sensor cards also reside.

DSC_2883_ss.jpg

  4411   Fri Mar 18 12:22:04 2011 kiwamuUpdateGreen LockingY arm plan for today

 Prior to the works on the Y end setup I propose to perform the temperature scan business like Koji and Suresh did before (see this entry).

This business will allow us to easily find a beatnote at 532nm after the installation on the Y end.

 I guess the right persons for this work are Bryan and Suresh.

Bryan will have a safety guidance from Steve in this after noon. So after that they can start working on it.

 

/* - - - coarse plan - - - */

* remove Alberto's laser from the AS table

* setup Alberto's laser on the PSL table

* put some stuff such as lenses, mirrors and etc. (Use the IR beam picked off after the doubling crystal for the main laser source)

* mode matching

* measurement

 

Which laser are we going to use,  Alberto's laser or MOPA laser ?

  4410   Fri Mar 18 11:29:36 2011 josephbUpdateCDSMinute trend issues

[Joe, Alex]

Steve pointed out to me today he couldn't get trends for his PEM slow channels like C1:PEM-count_full. 

I experimented a bit and found for long time requests (over 20 days), it would produce minute trends up to the current time, but only if they started far enough back.  So the data was being written, but something was causing a problem for dataviewer/NDS to find it.

On further investigation it looks to be some incorrect time stamps at several points in the last few months are causing the problems.  Basically when Alex and I made mistakes in the GPS time stamp settings for the frame builder (daqd) code, the wrong time got written for hours to the raw minute trend data files.

So Alex is going to be running a script to go through the roughly 180 gigabytes of affected trend data to write new files with the correct time stamps.  Once it done, we'll move the files over.  We'll probably lose a few hours worth of recent trend data, depending on how quickly the scripts run, but after which minute trends should work as they are supposed to.

  4409   Sun Mar 13 16:46:48 2011 josephbUpdateCDSETMY Sim work

4. The blue Output Filters  section has been changed to agree with the new filter of matrices row, column labeling.  My fault for not testing it and realizing it was broken.  The change was made in /opt/rtcds/caltech/c1/medm/master/C1SUS_DEFAULTNAME.adl and then ,/generate_master_screens.py was run, updating all the screens.

5.  I have swapped the logic for the sensor filter banks (ULSEN, URSEN, etc).  It now sends a "1" to the Binary Output board controlling the OSEM analog whitening when the FM1 filter is ON.  This has been done for all the suspensions (BS, ITMX,ITMY, SRM, PRM, MC1, MC2,MC3, ITMX, ITMY).

I am also updating the first sensor filter banks for the BS, ITMX, ITMY, SRM, PRM,MC1,MC2,MC3, called "3:30", to match the Y and X ends.

8.  I can't find any documentation on how to get a momentary button press to toggle states.  I could stick a filter bank in and use the on/off feature of that part, but that feels like a silly hack.  I've decided for the moment to split the TM offset button into 2, one for ON, one for OFF.  I'll put in on the list of things to have added to the RCG code (either a method, or documentation if it already exists). 

EDIT:  TM offset still doesn't work.  Will worry about it next week.

9.  Fixed a connection in SPY/SPX models where  the side senor path that was missing a constant to a modulo block.

Quote:

I did some work on the ETMY real and Sim.

  1. Set SUS coil gains to have the same quadropole arrangement as the magnets do (-1, 1, 1, -1) so that POS = POS instead of pringle.
  2. Set the Sim Magnet polarities to match this. These are the ETMY_CI filter banks.
  3. Found that the Xycom cable from the ETMY slow controls was unplugged at the Xycom side. This was preventing enabling the ETMY coil driver and so there was no real damping of the suspension going on. I plugged it in and checked that the mirror could now be moved.
  4. The C1SUS_ETMY master screen's BLUE output filter area is now mis-labeled. If you trust the screen you would set it up to drive the suspension incorrectly. This MUST be fixed along with all of the other misleading features of the screen.
  5. ETMY SUSSIDE filter bank had a 2048 Hz sample rate and was making the damping not work correctly. Fixed to 16384 Hz.
  6. 12 Hz, 4th order Cheby low pass added and turned on for the local damping filtering. This is not optimum, but its just there to give us some filtering without introducing some instability via phase lag around 3 Hz.
  7. ETMY OL beam re-aligned on ISCT-EX.
  8. TM Offset buttons not working on the main overview screen.

It seems like there is still a problem with the input whitening filters. I believe the Xycom logic is set such that the analog whitening of the OSEM signals is turned ON only when the FM1 is turned OFF. Joe has got to fix this (and elog it) so that we can damp the suspension correctly. For now, the damping of the ETMY and the SETMY require different servo gains and signs, probably because of this.

 

  4408   Sun Mar 13 04:00:53 2011 ranaUpdateCDSETMY Sim work

I did some work on the ETMY real and Sim.

  1. Set SUS coil gains to have the same quadropole arrangement as the magnets do (-1, 1, 1, -1) so that POS = POS instead of pringle.
  2. Set the Sim Magnet polarities to match this. These are the ETMY_CI filter banks.
  3. Found that the Xycom cable from the ETMY slow controls was unplugged at the Xycom side. This was preventing enabling the ETMY coil driver and so there was no real damping of the suspension going on. I plugged it in and checked that the mirror could now be moved.
  4. The C1SUS_ETMY master screen's BLUE output filter area is now mis-labeled. If you trust the screen you would set it up to drive the suspension incorrectly. This MUST be fixed along with all of the other misleading features of the screen.
  5. ETMY SUSSIDE filter bank had a 2048 Hz sample rate and was making the damping not work correctly. Fixed to 16384 Hz.
  6. 12 Hz, 4th order Cheby low pass added and turned on for the local damping filtering. This is not optimum, but its just there to give us some filtering without introducing some instability via phase lag around 3 Hz.
  7. ETMY OL beam re-aligned on ISCT-EX.
  8. TM Offset buttons not working on the main overview screen.

It seems like there is still a problem with the input whitening filters. I believe the Xycom logic is set such that the analog whitening of the OSEM signals is turned ON only when the FM1 is turned OFF. Joe has got to fix this (and elog it) so that we can damp the suspension correctly. For now, the damping of the ETMY and the SETMY require different servo gains and signs, probably because of this.

  4407   Sun Mar 13 00:00:58 2011 jzweizig, ranaConfigurationDAQNDS2 code change and restart

 John has changed the NDS2 code and restarted it on Mafalda. The issue is that it goes off the rails everytime the DAQD is restarted on FB because of filename convention war between GDS and CDS.

Until this is resolved, please make sure to restart the NDS2 process on Mafalda everytime you restart DAQD by doing this:

pkill -KILL nds2

/users/jzweizig/nds2-mafalda/start_nds2

  4406   Fri Mar 11 18:32:45 2011 josephb, Chris, JamieUpdateCDSDebugging simplant damping

The FM1 filter module change for XXSEN was propagated to the ETMX suspension.  This was a change from a 30,100:3 with a DC gain of 1 to a 3:30 which just compensates the hardware filter.

The good gains for the Sim damping were found to be:  100 for ETMX_SUSPOS, 0.1 ETMX_SUSPIT, and 0.1 ETMX_SUSYAW, ETMX_SUSSIDE is -70.  Gains much higher tended to saturate the simulated coils (i.e. hitting 10V limit) and then had to have the histories cleared for the RESPONSE matrix.

These seem to work to damp the real ETMX as well.

  4405   Fri Mar 11 16:30:42 2011 steveUpdateGeneralhigh speed servo unit is here

New Focus Servo Controller has just arrived. We have 25 days to evaluate this product.

It will have to be shipped back to the vendor on April 4, 2011 the latest in order to get full refund.

Attachment 1: P1070462.JPG
P1070462.JPG
  4404   Fri Mar 11 11:33:24 2011 josephbUpdateCDSFixed mistake in Matrix of Filter banks naming convention

While fixing up some medm screens and getting spectra of the simulated plant, I realized that the naming convention for the Matrices of Filter banks was backwards when compared to that of the normal matrices (and the rest of the world).  The naming was incorrectly column, row.

This has several ramifications:

1) I had to change the suspensions screens for the TO_COIL  output filters.

2) I had to change the filters for the suspension with regards to the TO_COIL output filters so they go in the correct filter banks.

3) Burt restores to times previous March 11th around noon, will put your TO_COIL output filters in a funny state that will need to be fixed.

4) The simplant RESPONSE filters had to be moved to the correct filter banks.

5) If you have some model I'm not aware of that uses the FiltMuxMatrix piece, it is going to correctly build now, but you're going to have to move filters you may have created with foton.

  4403   Thu Mar 10 21:45:34 2011 ranaUpdateGreen LockingIntensity stabilization loop using beatnote DC.

Ridiculous and hacky. Digital stabilization removed as well as the old "leave a pile of equipment on a stool" strategy.

We used a a BNC cable to send a pickoff of the beam before the recombination to the end via an SR560.

  4402   Thu Mar 10 17:03:48 2011 Larisa ThorneConfigurationElectronicscalculations for passive low pass filter on X arm

[Kiwamu, Larisa] 

 

We want to increase gain in the lower frequencies, so a circuit must be designed (a passive low pass filter). 

 

First, measurements were taken at the X arm for impedance and capacitance, which were 104.5kOhms and 84.7pF respectively. Kiwamu decided to make the circuit resemble a voltage divider for ease of calculation, such that Vout/Vin would be a ratio of some values of the equivalent circuit reactance values. After a few algebra mistakes, this Vout/Vin value was simplified in terms of the R, C measured and the R', C' that would be needed to complete the circuit. 

Since the measured C was very small and the measure R was fairly high, the simplified form allowed us to pick values of R' and C' that would make the critical frequency occur at 0.1Hz: set the R' resistance to 1MOhm and C' capacitance to 10uF, which would yield a gain ~1.

With these values a circuit we can start actually making the circuit.

  4401   Thu Mar 10 16:00:53 2011 Aidan, JoeUpdateGreen LockingIntensity stabilization loop using beatnote DC.

Aidan: Joe and I have added a channel that takes the DC output from the vertex beatnote PD and sends it, via RFM, to a DAC at the ETMX end. Immediately before the output is a filter C1GCX_AMP_CTRL. The output of the DAC is connected to the CURRENT LASER DIODE modulation input on the back of the Innolight driver. This will modulate the current by 0.1A/V input.

We should be able to modulate the green laser on the end now and stabilize the intensity of the amplitude on the beatnote PD at the vertex. (In this configuration, the ampltiude noise of the PSL laser will be injected onto the end laser - we may want to revisit that).

Joe's comments on model change:

I added a RFM connection at the output of the C1:GCV-XARM_BEAT_DC filter in the c1gcv model.  The RFM connection is called: C1:GCV-SCX_ETMX_AMP_CTRL.

This RFM connection goes to the c1scx model and into Kiwamu's GCX box, which uses top_names.  There's a filter inside called AMP_CTRL, so the full filter name is C1:GCX-AMP_CTRL.  The output then goes to the 7th DAC output.

Photos:

  1. NPRO CURRENT CTRL plugged into DAC channel 7
  2. You can actually see it's channel 7 in this image
  3. The other end plugged into the back of the Innolight driver
  4. Schematic of the setup
  5. Updated C1ALS_OVERVIEW MEDM screen (I don't know why the field in the background turned orange - maybe it's coming into a long dry summer?)

 

Quote:

There are 3 standard techniques to reduce this effect:

1) Stabilize the end laser by sensing the green light coming into the PSL before recombination and feeding back with SR560 (this is the only one that you should try at first).

 

The reason that I chose this PD is that, apparently, the green light coming from the cavity is clipped when it is picked off for its DC PD.

Attachment 1: P1000313.jpg
P1000313.jpg
Attachment 2: P1000314.jpg
P1000314.jpg
Attachment 3: P1000315.jpg
P1000315.jpg
Attachment 4: GREEN_ISS_LOOP.pdf
GREEN_ISS_LOOP.pdf
Attachment 5: Screenshot-C1ALS_OVERVIEW.adl.png
Screenshot-C1ALS_OVERVIEW.adl.png
  4400   Thu Mar 10 14:30:53 2011 ranaUpdateGreen LockingIntensity noise limits the beatnote sensitivity

There are 3 standard techniques to reduce this effect:

1) Stabilize the end laser by sensing the green light coming into the PSL before recombination and feeding back with SR560 (this is the only one that you should try at first).

2) Moving to the center of the MFD fringe via ETM steps.

3) Auto-alignment of the beam to the arm.

  4399   Thu Mar 10 14:29:05 2011 KojiUpdateGreen LockingIntensity noise limits the beatnote sensitivity

We can modify the freq divider circuit to make it a comparator.

Quote:

The next steps we should do are :

    - to investigate a cause of the intensity fluctuation
          * end green laser
          * suspensions' angular motions
          * doublecheck the RIN contribution if it's from the PSL or the X arm in the beatnote RFPD to make sure the RIN is dominated by the X arm transmitted light
  
    - to think about how to make the system insensitive to the intensity noise
          - bring the beat frequency to the zero cross point of the MFDs ?
          - PLL ?

Quote:

We are limited by the intensity noise of the X arm transmitted green light.

 

ELOG V3.1.3-